# A deep neural network approach to solve the Dirac equation

### v1

@@ Acknowledgements.
The authors thank Mario Centelles, Koji Hashimoto, and Hisashi Naito for the fruitful discussion. C.W. acknowledges the warm hospitality and computational resources of the RIKEN iTHEMS program. T.N. acknowledges the RIKEN Special Postdoctoral Researcher Program, the JSPS Grant-in-Aid for Research Activity Start-up under Grant No. JP22K20372, the JSPS Grant-in-Aid for Transformative Research Areas (A) under Grant No. JP23H04526, the JSPS Grant-in-Aid for Scientific Research (B) under Grant Nos. JP23H01845 and JP23K01845, the JSPS Grant-in-Aid for Scientific Research (C) under Grant No. JP23K03426, and the JSPS Grant-in-Aid for Early-Career Scientists under Grant No. JP24K17057. J.L. acknowledges the National Natural Science Foundation of China (Nos. 12475119 and 11675063), Scientific Research Project of Education Department of Jilin Province (No. JJKH20241242KJ), and Natural Science Foundation of Jilin Province (No. 20220101017JC). H.L. acknowledges the JSPS Grant-in-Aid for Scientific Research (S) under Grant No. JP20H05648 and the RIKEN Pioneering Project: Evolution of Matter in the Universe. The numerical calculations were performed on cluster computers at the RIKEN iTHEMS program.
@@ A.2 Usage of the trial wave function $f_{n\kappa}$
If $F_{n\kappa}$ is used as the DNN output, $G_{n\kappa}$ of a hydrogen atom diverges at the origin as shown in Fig. 14. The ground state wave functions of ${}^{208}\mathrm{Pb}$ are shown in Fig. 15. The diverge at the origin is smaller compared to the Coulomb results but still exists. This may because the gradient of $F_{n\kappa}$ of ${}^{208}\mathrm{Pb}$ at the origin is smaller compared with the hydrogen atom, which is easier for the DNN to optimize. So that the $G_{n\kappa}$ calculated by the $F_{n\kappa}$ from Eq. (11) is more accurate for ${}^{208}\mathrm{Pb}$ than the hydrogen atom.
@@ A.3 Minimizing without the inverse Hamiltonian
as the loss function, where the potential is the Coulomb potential as an example.
As shown in Fig. 16(a) for the non-fully connected DNN, it is observed that $\varepsilon_{\scriptstyle{\text{{{{D}}}}}}$ falls into the Dirac sea, which is the so-called variational collapse. For the fully connected DNN, as shown in Fig. 16(b) we discover that $\varepsilon_{\scriptstyle{\text{{{{D}}}}}}$ can converge at the ground state energy. This may because the local minimum of the ground state is too deep for the fully connected DNN so that the loss function is trapped.
@@
Figure 16: The energy of the Dirac Hamiltonian $H^{\prime}_{\scriptstyle{\text{{{{D}}}}}r}$ for a hydrogen atom with minimizing $\varepsilon_{\scriptstyle{\text{{{{D}}}}}}$ directly as a function of epochs. In the case of the non-fully connected DNN, the program doesn’t converge and falls into the Dirac sea. In the case of the fully connected DNN, the loss function converges at the ground state energy.

### v2

@@ Acknowledgements.
The authors thank Mario Centelles, Koji Hashimoto, and Hisashi Naito for the fruitful discussion. C.W. acknowledges the warm hospitality and computational resources of the RIKEN iTHEMS program. T.N. acknowledges the RIKEN Special Postdoctoral Researcher Program, the JSPS Grant-in-Aid for Research Activity Start-up under Grant No. JP22K20372, the JSPS Grant-in-Aid for Transformative Research Areas (A) under Grant No. JP23H04526 and JP25H01558, the JSPS Grant-in-Aid for Scientific Research (S) under Grant No. JP25H00402, the JSPS Grant-in-Aid for Scientific Research (B) under Grant Nos. JP23H01845, JP23K26538, and JP25K01003, the JSPS Grant-in-Aid for Scientific Research (C) under Grant No. JP23K03426, the JSPS Grant-in-Aid for Early-Career Scientists under Grant No. JP24K17057, and the JSPS Grant-in-Aid for JSPS Fellows under Grant No. JP25KJ0405. J.L. acknowledges the National Natural Science Foundation of China (Nos. 12475119 and 11675063), Scientific Research Project of Education Department of Jilin Province (No. JJKH20241242KJ). H.L. acknowledges the JSPS Grant-in-Aid for Scientific Research (S) under Grant No. JP20H05648 and the RIKEN Pioneering Project: Evolution of Matter in the Universe. The numerical calculations were performed on cluster computers at the RIKEN iTHEMS program.
@@ A.2 Usage of the trial wave function $f_{n\kappa}$
If $F_{n\kappa}$ is used as the DNN output, $G_{n\kappa}$ of a hydrogen atom diverges at the origin as shown in Fig. 14. The ground-state wave functions of ${}^{208}\mathrm{Pb}$ are shown in Fig. 15. The diverge at the origin is smaller compared to the Coulomb case but still exists. This may be because the gradient of $F_{n\kappa}$ of ${}^{208}\mathrm{Pb}$ at the origin is smaller compared with the hydrogen atom, which is easier for the DNN to optimize. So that the $G_{n\kappa}$ calculated by the $F_{n\kappa}$ from Eq. (11) is more accurate for ${}^{208}\mathrm{Pb}$ than the hydrogen atom.
@@ A.3 Minimizing without the inverse Hamiltonian
as the loss function, where the Coulomb potential is adopted as a test case.
As shown in Fig. 16(a) for the non-fully connected DNN, it is observed that $\varepsilon_{\scriptstyle{\text{{{{D}}}}}}$ falls into the Dirac sea, which is the so-called variational collapse. For the fully connected DNN, as shown in Fig. 16(b) we discover that $\varepsilon_{\scriptstyle{\text{{{{D}}}}}}$ can converge at the ground-state energy. This may because the local minimum of the ground state is too deep for the fully connected DNN so that the loss function is trapped.
@@
Figure 16: The energy of the Dirac Hamiltonian $H^{\prime}_{\scriptstyle{\text{{{{D}}}}}r}$ for a hydrogen atom with minimizing $\varepsilon_{\scriptstyle{\text{{{{D}}}}}}$ directly as a function of epochs. In the case of the non-fully connected DNN, the program doesn’t converge and falls into the Dirac sea. In the case of the fully connected DNN, the loss function converges at the ground-state energy.